Monthly Archives: February 2012

NMRs are back up

All the spectrometers are operational after the power outage this morning.

Posted in Announcements | Comments Off

HMQC/HSQC/HMBC not work on I500

HMQC, HSQC and HMBC experiments currently are not working well on the I500 due to a hardware problem. All other experiments, including routine 1H, 13C, COSY, NOESY, ROESY, etc. work with no problem.

Posted in Announcements | Comments Off